OpenStudy (anonymous):

a man walks a certain distance at a certain speed . had he walked 1/2 km/hr faster,he would have taken one hour less. but if he had walked 1km/hr slower, he would have taken 3 hours longer. the distance travelled by the man is:?

OpenStudy (anonymous):

assume speed be 'x' , distance 'd' , time be 't' we know distance equals product of speed and time taken we can make three equations and then solve x=d/t (x+0.5)=d/(t-1) (x-1)=d/(t+3) now we can solve for distance 'd'..